OpenStudy (anonymous):

The slope of the line whose equation is 2x + 5y + 1 = 0 is

OpenStudy (ivettef365):

you can solve this problem by isolating the y, then you have formula y = mx + b where m = slope

OpenStudy (anonymous):

help me solve it

OpenStudy (anonymous):

see ... 5y= -2x-1 y=-2x/5-1/5 this will be the new equation where m = -2/5 so you have a negative slope

OpenStudy (anonymous):

did you get this??

OpenStudy (anonymous):

@chelseac2014 ??

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

no problem :)